Two firms, Monolyte and NuPlant, are emitting toxic smoke into the local town. Monolyte emits 5 tons, each of which would cost $4000 to clean up, while NuPlant emits 8 tons, each of which would cost $3000 to clean. The town wants to reduce pollution to a maximum of 6 tons. For each of the following, indicate what the total cost of cleanup would be for the firms (note: the firms will only clean up as much as they are required to—they are not running a charity; the town isn’t interested in spending taxpayer money on this, either).

a. The town limits each firm to a maximum of 3 tons of pollution.

b. The town gives each firm 3 tradable pollution permits.
Answer #1

a)

In case pollution is limited to 3 tons in each case.

Abatement required by Monolyte =(5-3)=2 tons of pollution,

Abatement cost for Monolyte=2*4000=$8000

Abatement required by NuPlant=(8-3)=5 tons of pollution,

Abatement cost for Nuplant=5*3000=$15000

Total Cost of abatement=8000+15000=$23000

b)

It is clear that abatement cost is lower in case of NuPlant.

In case, Monolyte has 3 permits, it will be required to abate 2 tons of emission.

It is economical for Monolyte to buy 2 permits from NuPlant (at a price between $3000 and $4000).

After permits are traded, Nuplant will have 1 pemit left and thus will be required to abate 7 tons of emissions.

Abatement cost for Nuplant will be equal to total abatement cost

So, Total abatement cost=7*3000=$21000